OpenStudy (anonymous):

How do i solve for x in the equation 7^x+1 =27?

OpenStudy (anonymous):

\[7^{x+1} =27\]

OpenStudy (anonymous):

x = log(26)/log(7)

OpenStudy (anonymous):

if u need more explanation tell me

OpenStudy (anonymous):

actually yes it would help

OpenStudy (anonymous):

wait ,the one is an power or is a normal number ?

OpenStudy (anonymous):

7 to the x+1 power

OpenStudy (anonymous):

then it will be x = (log(27)-log(7))/log(7)

OpenStudy (anonymous):

use the logarithm rule

OpenStudy (anonymous):

am i just converting logs to expoential ?

OpenStudy (anonymous):

yes,u convert the exponetial to logs